### v1

# Training LLMs to Better Self-Debug and Explain Code

Abstract
In the domain of code generation, self-debugging is crucial. It allows LLMs to refine their generated code based on execution feedback. This is particularly important because generating correct solutions in one attempt proves challenging for complex tasks. Prior works on self-debugging mostly focus on prompting methods by providing LLMs with few-shot examples, which work poorly on small open-sourced LLMs. In this work, we propose a training framework that significantly improves self-debugging capability of LLMs. Intuitively, we observe that a chain of explanations on the wrong code followed by code refinement helps LLMs better analyze the wrong code and do refinement. We thus propose an automated pipeline to collect a high-quality dataset for code explanation and refinement by generating a number of explanations and refinement trajectories and filtering via execution verification. We perform supervised fine-tuning (SFT) and further reinforcement learning (RL) on both success and failure trajectories with a novel reward design considering code explanation and refinement quality. SFT improves the pass@1 by up to 15.92% and pass@10 by 9.30% over four benchmarks. RL training brings additional up to 3.54% improvement on pass@1 and 2.55% improvement on pass@10. The trained LLMs show iterative refinement ability, and can keep refining code continuously. Lastly, our human evaluation shows that the LLMs trained with our framework generate more useful code explanations and help developers better understand bugs in source code.
1 Introduction
Code generation has become a crucial research task to automatically generate source code based on natural language description [1, 2, 3, 4]. Although the recent Large Language Models (LLMs) have shown impressive capability in code generation, generating the correct code for a complex problem in single attempt is still challenging [5, 6, 7, 8, 9, 10, 11, 12]. This is expected because even for human developers, completing a hard programming problem might need multiple rounds of trial-and-error debugging. Self-debugging capability that allows LLMs to retrospect the incorrect code and make changes to resolve the implementation errors is becoming increasingly more important besides their code generation ability.
Existing works [13, 14] investigate off-the-shelf LLMs in the scale of Codex (code-davinci-002[1], GPT-3.5 and GPT-4, and show that these LLMs are able to self-debug the wrong code they generated via prompting methods in a pipeline of code generation and self-refinement as shown in Figure 1. The user first queries the LLM for a solution for the given programming task and the initial solution from the LLM is verified by executing them against the given unit tests. If the solution passes all the unit tests, it is considered correct. Otherwise, the user collects the unit test feedback and forms a new query to ask the LLM for a refinement. Such a process can iterate until the LLM generates a correct solution or reaches the maximum number of iterations. There are different prompt designs when asking for refinement [13]. Compared with directly asking for a refined solution (referred to as “Code Refinement” in the feedback block), asking LLMs to provide an explanation of the wrong solution and then refine it in a chain-of-thought manner (referred to as “Code Explanation and Refinement” in the feedback block) helps it to better understand the unit test feedback and increases the success rate of providing refined solutions (details in Appendix A.1).
Figure 1: Pipeline of letting LLM generate code and self-debug.
However, how to improve LLMs’ self-debugging capability remains under-explored, especially given the fact that open-sourced LLMs such as StarCoder [11] and CodeLlama [12] have limited self-refinement performance. For example, the StarCoder-15B model is only able to refine 4.43% wrong solutions for problems from the MBPP benchmark [3], in contrast, GPT-3.5-Turbo can refine 28.90% under the same setting (details in Appendix A.1). Such limited self-refinement ability motivates the need to better train LLMs to take feedback to explain and self-refine the wrong code. Although important, an essential challenge of training LLMs to explain and refine wrong code is the lack of training data, especially high-quality code explanation data. Previous work has explored Imitation learning from Language Feedback (ILF) [15], which trains LLMs with human-annotated explanation, yet, such an approach is not scalable and the LLMs also do not obtain the ability to explain code.
In this work, we propose an automated pipeline to collect a high-quality dataset for code explanation and refinement by generating a number of explanation and refinement trajectories and filtering via execution verification. Supervised fine-tuning (SFT) on the high-quality demonstration data boosts LLMs’ wrong code explanation and code refinement ability with a significant margin compared to existing prompting approaches. Furthermore, we propose reinforcement learning (RL) with a novel reward design that considers code explanation semantics and unit test passing rate, enabling LLMs to generate more useful code explanations and correct refinements. To sum up, this work makes the following contributions:
  • We propose a scalable framework consisting of automated data collection, data verification, supervised fine-tuning, and reinforcement learning with new reward designs, to improve LLMs’ self-debugging ability and generate more helpful code explanations and correct refinements.
  • We perform supervised fine-tuning on three LLMs (StarCoder-15B, CodeLlama-7B, and CodeLlama-13B) with the collected code refinement and wrong code explanation data. Our extensive experiments demonstrate that it significantly improves their wrong code explanation and code refinement ability, with the pass@1 improved by up to 15.92% and the pass@10 improved by up to 9.30% on four benchmarks.
  • We train the fine-tuned LLMs using a reinforcement learning algorithm with a novel reward design considering unit test execution and wrong code explanation semantics, and further improve the pass@1 by up to 3.54% and pass@10 by up to 2.55% over SFT on four benchmarks.
  • We conduct comprehensive ablation and analytical experiments as well as a human evaluation to show the generalizability and effectiveness of our approach.
2 Approach
In this section, we present our proposed automated pipeline, which includes the collection of high-quality code explanation and refinement data, and the training methods. Figure 2 shows the overview of our approach. We first collect a code explanation and refinement dataset by querying from pretrained or instruct models and verifying its responses with execution feedback to filter and obtain high-quality explanation and refinement data (steps 1 and 2 in Figure 2, Section 2.1). Then the high-quality dataset is used for supervised fine-tuning (step 3 in Figure 2, Section 2.2), which significantly improves the model’s performance in explaining the bug and refining the code. Reinforcement learning with execution feedback is used to further guide the model to generate higher quality response and boost the model performance (step 4 in Figure 2, Section 2.3).
Figure 2: Overview of our approach.
2.1 Data collection and verification
We use MBPP [3] (only use the 374 problems in the training set during training), APPS [4] (only use the 5,000 problems in the training set) and CodeContests [2] as our base training datasets, which contain programming problems and solutions collected from various platforms. While they are helpful for training LLMs for code generation, they neither contain enough wrong solutions nor the explanation and refinement of them. To collect more wrong solutions, we prompt the pre-trained LLMs (i.e., StarCoder and CodeLlama) accordingly with 3-shot examples to sample 20 solutions (temperature set to 1.0) per problem from MBPP’s training set, APPS’s training set, and CodeContests. We then run these generations against test cases to select the wrong solutions that fail any test cases.
Table 1 shows the number of correct (passing all the unit tests) and wrong (failing any unit test) solutions sampled for each dataset. For each wrong solution, we need an explanation of the wrong code and a correct refinement to build the code explanation and refinement dataset. We prompt pre-trained or instruction-LLMs with the problem description, wrong solution, and execution feedback (either error message or failed test case) to ask for an explanation and refinement. In this paper, we experimented with GPT-3.5-Turbo and CodeLlama-34B. We take GPT-3.5-Turbo as the example LLM in this section, and an example with it is shown in Appendix A.2. We study the generalization of this data collection with different LLMs in Section 4.3.
Table 1: Number of unique, correct, wrong solutions sampled from pre-trained LLMs, as well as the number of correct refinement generated by GPT-3.5-Turbo and its refinement rate on each dataset.
Dataset #Unique Solutions #Correct Solutions #Wrong Solutions #Correct Refinement #Refinement Rate
MBPP Training (374) 9,500 4,706 4,794 2,203 45.95%
APPS Training (5,000) 44,108 27,736 16,372 6,419 39.21%
CodeContest (6,627) 51,134 31,520 19,614 5,113 26.07%
As LLMs may provide wrong explanations or refinements, we cannot blindly take them as training data. Thus, we verify the refinements by running them against the test cases again, and only those passing all the test cases are considered correct refinements. For explanation, we consider the explanations along with the correct refinements as correct. Overall, for example with GPT-3.5-Turbo, we get 13,735 correct explanations and refinements: 2,203 for MBPP, 6,419 for APPS, and 5,113 for CodeContests. This verification step is crucial to guarantee the quality of the automatically collected code explanation and refinement dataset.
2.2 Supervised fine-tuning
We form the fine-tuning data in an instruction-following format similar to StarChat [16], where the user input is enclosed by <|user|> and <|end|>, while LLM’s answer is enclosed by <|assistant|> and <|end|> in the chat history. Moreover, to alleviate the limited amount of data, we augment the fine-tuning data by using two different instructions: providing the task description, the initial wrong code, and execution feedback, asking for (1) a refinement directly, or (2) an explanation of the wrong code and then a refinement in a chain-of-thought manner. Examples are given in Appendix A.2
During supervised fine-tuning, although we include the wrong solutions as LLM’s initial answer in the chat history, we do not calculate the loss for this part since we do not want the LLM to intentionally generate those wrong solutions. They are just provided as context for code explanation and refinement if the LLM indeed makes mistakes in real use cases.
2.3 Reinforcement learning
Reinforcement learning is widely used to further improve the quality of LLM’s generated outputs [17, 18, 19, 20]. Through the RL framework, the LLM is optimized by using an algorithm to update the weights using both success and failure trajectories and maximize the rewards of its outputs. To train the fine-tuned LLMs to generate better code explanations and more correct code refinements, we design the rewards considering both parts.
2.3.1 Refinement score
To train LLM to refine code, the correctness of the refinement is the main goal, which can be measured by its code similarity to the ground truth, as well as the execution result. We use CodeBLEU score as metrics for code similarity and unit test passing rate as metrics for execution results.
Given a wrong solution $w$, the set of correct and wrong (failed) refinements are notated by $R^{w}_{c}$ and $R^{w}_{w}$. For any refinement $r$, we calculate its CodeBLEU score and the unit test passing rate as follows:
$S_{cb}(r)$ $=\frac{1}{|R^{w}_{c}|}\sum_{r_{c}\in R^{w}_{c}}\operatorname{CodeBLEU}(r,r_{c});\quad$ $S_{ut}(r)$ $=\frac{|T_{p}(r)|}{|T|}$
$S_{cb}$ is the average CodeBLEU score between a given refinement and all the correct refinements. $S_{ut}$ is the fraction of the number of passed unit test cases ($T_{p}$) when running the refined code $r$, over the total number of unit test cases ($T$) provided for this problem in the dataset.
In Figure 3, the x-axis is the scores of certain metrics, and the y-axis is the number of training data with a certain score (same for other figures in Figure 3). Thus, Figure 3 (a) shows the frequency distribution of each score of $S_{cb}$, with blue bars referring to training data with correct refinements, and orange bars referring to that with wrong refinements. The distribution of $S_{ut}$ is shown in Figure 3 (b) where the correct refinements definitely pass all the test cases and can be separated from the wrong ones.
Figure 3: The CodeBLEU scores, unit test cases passing rate, sentiment similarity of wrong code explanations, final refinement code reward, and the explanation reward of the training data.
2.3.2 Explanation score
In our dataset, there are wrong code explanations along with code refinement, whose quality may not be perfectly reflected by the code quality. A correct code explanation may also be followed by incorrect refinement, thus, it is necessary to consider the explanation in the reward. The code explanations followed by correct refinements are treated as ground truth, notated by $E^{w}_{c}$. We calculate the average sentiment similarity [21, 22] between the explanation embedding $e$ and corresponding embeddings in ground truth as
$\footnotesize S_{ex}(e)=\frac{1}{|E^{w}_{c}|}\sum_{e_{c}\in E^{w}_{c}}\operatorname{CosSim}(\operatorname{RoBERTa}(e),\operatorname{RoBERTa}(e_{c}))$
The distribution of $S_{ex}(e)$ is shown in Figure 3 (c).
2.3.3 Reward design
Given a pair of explanation and refinement, i.e., $(e,r)$, the reward of the generated code refinement and the code explanation is designed as:
$\mathcal{R}(r)$ $=5\cdot(S_{cb}(r)+S_{ut}(r))-5;\quad$ $\mathcal{R}(e)$ $=\frac{50\cdot S_{ex}(e)-35}{3}$
This code reward $\mathcal{R}(r)$ is the average of CodeBLEU score and the unit test passing rate, and since both $S_{cb}(r)$ and $S_{ut}(r)$ are scored in the range of [0, 1], this equation makes the reward of code refinement in the range of [-5, 5]. Figure 3 (d) shows the distribution of the code refinement reward on the training dataset, which mitigates the overlap issue between correct and wrong outputs with CodeBLEU score alone as illustrated in Figure 3 (a). It also makes the reward distribution continuous, addressing the discreteness problem of only using unit test passing rate.
For the design of explanation reward $\mathcal{R}(e)$, We observe from Figure 3 (c) that the explanation sentiment similarities of the training data mostly lie in the range of [0.4, 1.0], thus, we project the range of [0.4, 1.0] to [-5, 5] and treat 0.7 as the borderline (projected to 0 correspondingly) of good or bad explanations. Figure 3 (e) shows the distribution of the wrong code explanation reward on the training dataset. The distribution shows that there could be a good or correct code explanation followed by a wrong code refinement, where assigning a high or low reward to the entire output is not reasonable. This leads to our PPO [23] algorithm with code refinement and explanation rewards considered separately. Due to space limit, the PPO algorithm is shown in Appendix A.3.
3 Experimental setup
For supervised fine-tuning, we fine-tune three LLMs (StarCoder-15B, CodeLlama-7B, and CodeLlama-13B) using the correct initial solutions and correct refinements collected from the MBPP training set, APPS training set, and CodeContests. The model is fine-tuned for two epochs, using a batch size of 128. The optimizer is AdamW [24] with learning rate set to $2e^{-5}$. The learning rate is adjusted using a warmup of 500 steps and then decayed following a cosine scheduler.
We further train supervised fine-tuned LLMs with reinforcement learning using the PPO algorithm. The reinforcement learning training data is all the initial solutions and collected refinement on the MBPP and APPS training set. The learning rate is $2e^{-6}$, and the batch size is set to 64. We implement reinforcement learning training based on the TRL [25] library. Both the supervised fine-tuning and reinforcement learning are conducted on 8 NVIDIA A100 GPUs, each with 40GB of memory.
4 Result
To evaluate the effectiveness of our training approach, we study the following research questions (RQs) regarding code generation and code refinement capability, iterative refinement ability, approach generalizability, and the quality of the generated code explanations.
4.1 RQ1: Code generation and refinement capability
We evaluate the models trained with our approach for their code explanation and refinement ability using four benchmarks: MBPP [3], HumanEval [1], MBPP+ [26], and HumanEval+ [26]. We use pass@k [1] and success refinement rate as the evaluation metric. For the generation of the initial solutions, the models sample 100 solutions per task in the benchmarks (temperature set to 0.8), which are run against the provided test cases. For every incorrect solution that fails any test case, we let the models sample one refinement (and one explanation).
4.1.1 Pass@k
Table 2 shows the pass@k result on four benchmarks separately. In general, fine-tuning LLMs with our collected code explanation and refinement data significantly improves the pass@1 and pass@10 of all three backbones. For both StarCoder-15B and CodeLlama-13B, RL achieves the best pass@1 and pass@10 (bolded) on all four benchmarks. For CodeLlama-7B, RL achieves the best performance under seven settings except that SFT gets the highest pass@1 on the MBPP benchmark.
Table 2: Pass@k of initial and refined solutions on four benchmarks. Each LLM’s best performance on every benchmark is bolded.
Models Approaches MBPP Humanval MBPP+ HumanEval+
pass@1 pass@10 pass@1 pass@10 pass@1 pass@10 pass@1 pass@10
Init. 37.70 69.60 30.34 63.21 34.90 60.85 26.16 56.77
Prompt. Refine 40.64 71.04 34.05 65.42 40.48 66.59 30.73 61.11
Expl. + Refine 40.37 71.60 33.96 67.20 39.23 64.67 30.09 60.72
Init. 47.41 70.27 35.05 66.93 43.28 62.48 30.01 59.77
StarCoder-15B SFT Refine 56.66 75.78 47.32 77.06 53.53 71.55 43.16 71.74
Expl. + Refine 57.11 76.70 47.37 77.16 53.83 72.19 43.54 72.61
Init. 48.62 70.68 39.45 70.16 44.94 63.92 35.05 64.41
RL Refine 58.00 78.12 52.25 80.80 54.13 71.71 46.11 74.07
Expl. + Refine 58.19 77.96 51.67 80.79 54.29 71.93 46.26 74.24
Init. 38.21 67.24 34.27 69.60 37.18 61.23 27.40 60.81
Prompt. Refine 43.98 71.94 39.20 74.14 42.97 66.89 31.84 65.08
Expl. + Refine 43.42 72.09 40.13 74.95 42.46 67.41 32.49 66.58
Init. 48.87 70.89 36.99 69.95 42.97 62.69 30.76 62.52
CodeLlama-7B SFT Refine 58.07 77.34 52.65 80.71 51.64 71.04 46.61 74.43
Expl. + Refine 57.98 77.92 52.98 82.22 51.55 70.94 47.62 75.54
Init. 46.54 71.54 39.38 71.84 41.46 63.68 33.95 65.98
RL Refine 57.35 78.12 54.41 82.55 52.11 71.88 48.73 77.17
Expl. + Refine 57.92 78.97 55.84 84.14 52.90 71.80 50.04 78.25
Init. 42.88 70.85 37.11 74.69 38.93 62.26 30.15 66.27
Prompt. Refine 49.68 75.85 45.78 81.07 46.22 70.14 37.62 72.68
Expl. + Refine 49.97 76.39 45.90 81.18 45.77 70.48 38.36 73.84
Init. 52.43 73.66 41.65 73.61 43.67 62.63 35.29 68.49
CodeLlama-13B SFT Refine 61.78 79.96 58.41 83.47 55.58 73.41 51.35 77.84
Expl. + Refine 61.59 80.21 57.76 84.57 54.59 72.15 51.32 78.84
Init. 51.19 73.16 45.45 74.79 45.49 62.81 39.27 69.29
RL Refine 61.98 79.95 61.71 84.58 57.89 73.48 56.68 80.89
Expl. + Refine 61.63 80.27 61.66 86.23 56.62 72.04 56.57 81.77
For simpler comparison, Table 3 shows the overall pass@k on MBPP and HumanEval, as well as the improvement brought by SFT and RL. The improvement by SFT is compared to prompting, and the improvement by RL is compared to SFT.
On the overall of MBPP and HumanEval, SFT improves StarCoder-15B by a 15.34% higher pass@1 and a 6.58% higher pass@10 after direct refinement. When asking for code explanation and refinement in the chain of thought, the SFT StarCoder-15B obtains a 15.92% and 6.30% improvement on pass@1 and pass@10. RL training improves SFT by another 2.43% pass@1 and 2.55% pass@10 when directly generating refinement, and improves by another 1.87% pass@1 and 1.85% pass@10 when generating code explanation and refinement. Similar improvement by SFT and RL can also be observed on CodeLlama-7B and 13B backbones.
Table 3: Overall pass@k on MBPP & HumanEval and MBPP+ & HumanEval+. ( ) shows the improvement. SFT’s improvement is compared to prompting, and RL’s is compared to SFT.
StarCoder MBPP & HumanEval MBPP+ & HumanEval+
15B SFT RL SFT RL
Refine Expl. + Ref. Refine Expl. + Ref. Refine Expl. + Ref. Refine Expl. + Ref.
pass@1 54.35 (+15.34) 54.71 (+15.92) 56.78 (+2.43) 56.58 (+1.87) 49.31 (+12.80) 49.64 (+14.13) 50.87 (+1.56) 51.02 (+1.38)
pass@10 76.23 (+6.58) 76.81 (+6.30) 78.78 (+2.55) 78.66 (+1.85) 71.63 (7.27) 72.36 (+9.30) 72.67 (+1.04) 72.87 (+0.51)
CodeLlama MBPP & HumanEval MBPP+ & HumanEval+
7B SFT RL SFT RL
Refine Expl. + Ref. Refine Expl. + Ref. Refine Expl. + Ref. Refine Expl. + Ref.
pass@1 57.21 (+14.41) 56.75 (+14.14) 56.62 (-0.60) 57.41 (+0.66) 49.59 (+11.15) 49.95 (+11.55) 50.73 (+1.14) 51.74 (+1.79)
pass@10 78.17 (+5.69) 78.98 (+6.18) 79.21 (+1.04) 80.25 (+1.27) 72.42 (+6.27) 71.81 (+4.74) 74.03 (+1.61) 74.42 (+2.60)
CodeLlama MBPP & HumanEval MBPP+ & HumanEval+
13B SFT RL SFT RL
Refine Expl. + Ref. Refine Expl. + Ref. Refine Expl. + Ref. Refine Expl. + Ref.
pass@1 60.95 (+12.23) 60.64 (+11.68) 61.91 (+0.96) 61.64 (+1.00) 53.86 (+11.14) 53.26 (+10.51) 57.40 (+3.54) 56.60 (+3.34)
pass@10 80.83 (+3.69) 81.29 (+3.72) 81.09 (+0.26) 81.74 (+0.45) 75.21 (+4.04) 74.87 (+3.02) 76.50 (+1.29) 76.00 (+1.13)
On the overall of MBPP+ and HumanEval+ benchmarks, where each problem has 35$\times$ and 80$\times$ more rigorous test cases [26], we observe larger improvements from RL training on CodeLlama backbones. CodeLlama-7B achieves a 1.79% higher pass@1 and a 2.60% higher pass@10 on the refined solutions with code explanation. CodeLlama-13B achieves a 3.54% higher pass@1 and a 1.29% higher pass@10 on the directly generated refinements. This shows that RL training enables LLMs to generate or refine more robust solutions that can pass rigorous test cases. More experiments and examples of case studies can be found in Appendix A.4.1,  A.5.1,  A.5.2,  A.5.3.
4.1.2 Success refinement rate
Table 4 lists the success refinement rate generated by each backbone using different approaches. The result is averaged over four benchmarks. For StarCoder-15B, prompting is only able to correctly refine 6.41 – 6.90% of the wrong initial solutions. After being fine-tuned with the collected code explanation and refinement data, the supervised fine-tuned StarCoder-15B refines 16.27 – 16.56% of the wrong solutions with a significant improvement of 9.37 – 10.15%. RL training further increases the refinement rate by 1.03 – 1.23% over SFT.
Table 4: Success refinement rate of different approaches over four benchmarks.
Models Refine (%) Explain + Refine (%)
Prompt. SFT RL Prompt. SFT RL
StarCoder-15B 6.90 16.27 (+9.37) 17.50 (+1.23) 6.41 16.56 (+10.15) 17.59 (+1.03)
CodeLlama-7B 8.65 18.14 (+9.49) 19.95 (+1.81) 8.10 17.60 (+9.50) 20.84 (+3.24)
CodeLlama-13B 11.64 18.96 (+7.32) 22.58 (+3.62) 11.97 20.06 (+8.09) 23.50 (+3.44)
The improvement on CodeLlama-7B and 13B backbones is consistent with that on StarCoder-15B, where RL training eventually achieves the highest success refinement rate with a considerable boost of 1.81 – 3.62%.
4.2 RQ2: Iterative refinement ability
LLMs can self-debug iteratively until they find the correct solutions. Figure 4 shows the overall pass@k of CodeLlama-7B on four benchmarks after at most three rounds of refinements. To simplify the figure, we plot the higher pass@k from "Refine" and "Expl. + Refine" in each round of refinement of each LLM. More results of iterative refinement are in Appendix A.4.2.
Figure 4: Pass@k of prompting, SFT, and RL CodeLlama-7B after three iterations of refinements.
Both SFT and RL consistently outperform prompting in all three rounds of refinement. Even with three rounds of refinements, the prompting approach cannot achieve the same pass@k as SFT (e.g., 47.53% vs. 56.75% in Figure 4 (a)) after the first round of refinement. These results show that LLMs trained with our pipeline have iterative refinement ability and can achieve increasingly higher pass@k with more iterations.
4.3 RQ3: Generalizability of approach
To show the generalizability of our framework, i.e., especially that the data collection is independent of GPT-3.5-Turbo, we use CodeLlama-34B as an alternative to GPT-3.5-Turbo for data collection. Since CodeLlama-34B is a pre-trained model, we use few-shot examples in the prompt to make it generate the desired wrong code explanation and refinement. The rest remains the same.
Table 5: Pass@k on MBPP and HumanEval by LLMs trained with CodeLlama-34B’s data.
Models Approaches MBPP Humanval MBPP+ HumanEval+
pass@1 pass@10 pass@1 pass@10 pass@1 pass@10 pass@1 pass@10
CodeLlama-7B Init. 47.01 71.18 39.24 70.00 41.64 61.95 33.07 63.58
SFT Refine 54.97 76.52 47.03 77.96 48.71 68.28 41.11 71.46
Expl. + Refine 55.05 76.63 47.34 78.11 49.15 68.35 41.49 71.76
Init. 46.45 70.41 39.63 71.02 42.34 60.84 34.15 63.19
RL Refine 54.81 76.56 48.28 80.06 53.06 70.17 47.40 73.81
Expl. + Refine 55.32 76.74 48.52 79.14 53.17 69.36 48.59 75.80
Table 6: Overall pass@k on MBPP & HumanEval and MBPP+ & HumanEval+, trained with CodeLlama-34B’s data.
CodeLlama MBPP & HumanEval MBPP+ & HumanEval+
7B SFT RL SFT RL
Refine Expl. + Ref. Refine Expl. + Ref. Refine Expl. + Ref. Refine Expl. + Ref.
pass@1 50.01 (+7.21) 53.15 (+10.54) 53.20 (+0.05) 53.64 (+0.49) 45.62 (+7.18) 46.03 (+7.63) 50.76 (+5.14) 51.31 (+5.28)
pass@10 76.88 (+4.40) 77.00 (+4.20) 77.42 (+0.42) 77.33 (+0.33) 69.57 (+3.42) 69.74 (+2.67) 71.65 (+2.08) 71.98 (+2.24)
Table 5 shows the pass@k results of CodeLlama-7B trained with data collected from CodeLlama-34B and Table 6 shows the overall for simpler comparison. SFT obtains smaller (around 1%-3% less compared to training with GPT-3.5-Turbo’s data) but still notable improvements on the overall pass@1 and pass@10. Furthermore, we observe the same finding that RL training improves more of the performance on MBPP+ and HumanEval+ by 5.28% of pass@1 and 2.24% of pass10. This shows the generalizability of our pipeline and also suggests that collecting data from a stronger LLM benefits the training more in our pipeline. More results are shown in Appendix A.4.3.
4.4 RQ4: Quality of generated explanation
We evaluate if the explanation for wrong code is useful for developers to understand their bugs. We randomly sample 50 problems with initial wrong solutions from the MBPP and HumanEval benchmarks and query different LLMs to generate explanations for the wrong code. We score each code explanation on a scale of 1 – 5 considering their correctness and helpfulness, with 1 being totally incorrect or misleading, and 5 being not only correct but also providing a detailed hint of how to fix the wrong code.
We perform scoring using both GPT-4 and human developers as judges. When using GPT-4 as a rater, we follow previous work [27] to prompt it to score each explanation, and the result is shown in Table 7. Both SFT and RL improve the quality of explanation over the prompting approach. Overall, GPT-4 gives higher scores to the LLMs trained with our approach than the prompting approach. Compared with GPT-3.5-Turbo explanations, the gap is significantly reduced after the training.
Table 7: Average scores of code explanations rated by GPT-4 and developers. SC for StarCoder and CL for CodeLlama. “-” refers to not applied.
Raters Prompting SFT RL Prompting SFT RL Prompting SFT RL Prompting
SC-15B SC-15B SC-15B CL-7B CL-7B CL-7B CL-13B CL-13B CL-13B GPT-3.5-Turbo
GPT-4 1.90 2.92 3.18 2.00 3.00 3.10 2.54 2.62 3.24 3.60
Developers 1.73 2.60 2.76 - - - - - - 3.21
Due to the effort required by human rating, we only let developers rate the StarCoder models’ and GPT-3.5-Turbo’s generations and each explanation is rated as the average of scores given by two developers. The human rating is consistent with GPT-4 indicating that SFT improves the explanation quality over prompting and RL training takes code explanation semantics into consideration of rewards, thus bringing continuous improvement to the explanation generation. Rubrics and examples of human rating are shown in Appendix A.6.
5 Limitation
One potential limitation of our study is the reliance on specific large language models (LLMs) from which we collect code explanation and refinement data. Our automated framework is designed to be independent of any specific LLM, and for this study, we use GPT-3.5 and CodeLlama-34B to collect training data, and both bring significant improvement through SFT and RL. However, for future work, it would be interesting to explore the use of other LLMs, including smaller models or a mix of diverse LLMs, to gather explanation and refinement data.
Additionally, our current experiments only use two types of prompts for enabling LLMs to self-debug: one that directly asks for refinement and another that first asks for an explanation of the wrong code followed by refinement. While these prompt designs have shown effectiveness, there might be better prompt strategies for self-debugging that we have not explored due to resource constraints. Exploring a broader range of prompt designs could potentially enhance the performance of our framework. Nonetheless, our proposed training framework is flexible and should be generalizable to different types of data and prompts, paving the way for future innovations in this area.
6 Related work
6.1 Large Language Models for Code
Large language models (LLMs) have been widely explored across a variety of code-related tasks, including code generation [5, 8, 6, 7, 10, 9, 12, 28, 29, 30, 31, 32, 33, 34], bug fixing [35, 36, 37, 38], program testing [39, 40] and fuzzing [41] and so on. These models have demonstrated impressive capabilities in these domains, largely due to their strong understanding and generation abilities acquired through extensive pre-training on vast datasets. This pre-training allows them to recognize patterns, understand context, and generate coherent and contextually relevant code snippets.
However, most existing works in this area focus primarily on improving LLMs to provide the expected output in a single round of generation. The emphasis has been on enhancing the initial output quality, minimizing the need for further modifications or iterations. This one-shot generation approach, while useful, overlooks the potential of iterative refinement, which is a crucial aspect of real-world programming where initial drafts often require multiple rounds of revision and debugging.
6.2 Self-Debugging
Existing techniques have studied the possibility of using LLMs to refine their generations. Yet, most techniques are prompting LLMs with execution results [13, 14, 42, 43, 44, 45, 46, 47] for the refinement. Such prompting approaches bring limited improvement to smaller open-sourced LLMs compared to GPT-3.5. Other techniques train LLMs to self-debug. ILF [48] uses human-annotated feedback information and thus is unscalable, CYCLE [49] and Self-Edit [50] use SFT to fine-tune LLM to generate the refinement only based on the unit test execution feedback.
This work has three differences compared with others that train LLMs: (1) we train LLMs to generate code explanation followed by refinement, which provides additional information to users, (2) we do not require human-annotated training data but propose a scalable pipeline to automatically collect and verify data from another LLM, (3) we design novel reward functions in the RL training stage, considering both the code and explanation quality, which beings extra improvement.
7 Conclusion
This work shows the necessity of training open-sourced LLMs to self-debug and proposes a scalable framework consisting of automated data collection, data verification, supervised fine-tuning, and reinforcement learning with new reward designs, to improve LLMs’ self-debugging ability. The data collection is independent of LLMs. Our experiments using either GPT-3.5-Turbo or CodeLlama-34B achieve great improvements. The data verification ensures the high quality of the collected code explanation and refinement data. SFT on our collected data brings a significant boost to LLMs’ self-debugging ability, resulting in up to 15.92% higher pass@1, 9.30% higher pass@10, and 10.15% higher success refinement rate. RL training using our novel reward design further brings up to 3.54% higher pass@1, 2.55% higher pass@10, and 3.62% higher success refinement rate. Comprehensive analytical experiments validate the generalizability of our approach and illustrate the iterative refinement ability of the trained LLMs. The human evaluation indicates that the explanations generated by the LLMs trained with our approach are of higher quality and can effectively serve developers for debugging purposes.
Appendix A Appendix
A.1 Limited refinement ability of open-sourced LLMs
We test two prompts of querying LLMs for code refinement, letting LLMs generate the code refinement only (Figure 5 (3.1)), or letting LLMs first explain why the code is wrong and then generate the refinement (Figure 5 (3.2)).
Figure 5: Two different prompts to ask LLM to self-refine: directly asking for refinement (left), asking for an explanation of the wrong code, and then refining in chain-of-thought (right).
Table 8 shows the success rate of refinement generated by StarCoder-15B, CodeLlama-7B, CodeLlama-13B, and GPT-3.5 using greedy decoding. Mostly, when LLMs explain the wrong code first, they are more likely to provide the correct refinement. Yet, all the three open-source LLMs we test, even with the best efforts, show limited code refinement ability. For instance, StarCoder-15B is only able to refine 4.43 – 5.10% of the wrong code, which motivates the training of open-sourced LLMs to obtain code explanation and refinement ability.
Table 8: The success rate of self-refinement using greedy decoding.
Models MBPP HumanEval
Refine Expl. + Refine Refine Expl. + Refine
StarCoder-15B 2.58% 4.43% 2.04% 5.10%
CodeLlama-7B 7.42% 6.71% 5.32% 8.52%
CodeLlama-13B 9.80% 10.20% 5.88% 8.24%
GPT-3.5 26.24% 28.90% 32.32% 33.33%
A.2 Data collection and data format for training
Figure 6 shows an example of the prompt we used to collect code explanation and refinement training data from GPT-3.5 (same for collecting data from CodeLlama-34B). We use a temperature of 0.8 to let GPT-3.5 sample 10 code explanations and refinements per prompt.
Figure 6: Prompt used for code explanation and refinement data collection.
The data collected contains both the code explanation and refinement, and we construct two formats of instruction-tuning data for SFT. Figure 7 (left) shows the instruction asking LLMs for code refinement only, where the instruction contains the task description, wrong initial solution, and execution feedback. LLMs are trained to generate the output refinement given the instruction. Figure 7 (right) shows another instruction asking LLMs for both code explanation and refinement. We include both types of instruction data in the training, which increases the training data diversity and enhances LLMs’ robustness against different prompts.
Figure 7: Two types of instruction tuning data used in SFT and RL.
A.3 PPO algorithm with separate rewards in RL
We modify the standard PPO algorithm to optimize the explanation and refinement generation based on their reward separately. Given an LLM (already supervise-fine-tuned) $\pi$ as the policy model, a prompt $x$, a generation $y$ consists of a code explanation and a refinement:
$\footnotesize y=[e,r]=\{y_{1},y_{2},\ldots,y_{|e|},y_{|e|+1},y_{|e|+2},\ldots,y_{|e|+|r|}\}$
$\{y_{i}\}_{i=1}^{|e|}$ is the explanation of the wrong code, and $\{y_{i}\}_{i=|e|}^{|e|+|r|}$ is the refinement. We then define the advantage $A$ in the PPO algorithm as:
$A_{t}=\delta_{t}+\gamma\delta_{t+1}+\ldots+\gamma^{\mathcal{T}-t}\delta_{\mathcal{T}}\quad,$ $\delta_{t}=\mathcal{R}_{t}-V(y_{<t},x,\pi)+\gamma V(y_{<t+1},x,\pi)$
$A_{t}$ is the advantage at decoding timestamp $t$, $\mathcal{T}=|e|+|r|$ is the total length of generation output, and $\gamma$ is the discount rate (a hyper-parameter set to 0.99 in our experiment). $V(y_{<t},x,\pi)$ is the state value at generation step $t$ given input $x$, which is learned and calculated by a linear layer on top of the policy model $\pi$. $\mathcal{R}_{t}$ is the reward at decoding timestamp $t$, which is calculated as follow:
$\footnotesize\mathcal{R}_{t}=\left\{\begin{aligned} &\mathcal{R}(r)-\operatorname{KL}_{t}(\pi,\pi^{\prime})\qquad\qquad\qquad\qquad,\;t=\mathcal{T}\\ &\mathcal{R}(e)-\operatorname{KL}_{t}(\pi,\pi^{\prime})\qquad\qquad\qquad\qquad,\;t=|e|\\ &\operatorname{KL}_{t}(\pi,\pi^{\prime})\approx\operatorname{log}\frac{\operatorname{P}(y_{t}|y_{t-1},x,\pi)}{\operatorname{P}(y_{t}|y_{t-1},x,\pi^{\prime})}\;\;\;\;,\;\text{otherwise}\end{aligned}\right.$
where $\operatorname{KL}$ is the Kullback–Leibler divergence [51] between the action distribution given by the updated policy model $\pi$ and the old policy model $\pi^{\prime}$ before the update.
Instead of assigning both the code and explanation rewards to the entire output, we separate them by only assigning the explanation reward to the explanation portion. This avoids the issue that a low reward is assigned to a correct explanation followed by a wrong refinement, and the LLM learns to keep the correctly generated explanation part and focus on improving the incorrect refinement portion. For data where only code refinement but no explanation is generated, we use the standard design to assign the code reward to the code refinement.
Following existing works, the loss of the PPO algorithm training is:
$\footnotesize L=-\mathbb{E}\left[\sum_{t=1}^{\mathcal{T}}\frac{\operatorname{log}\operatorname{P}(y_{t}|y_{<t},x,\pi)}{\operatorname{log}\operatorname{P}(y_{t}|y_{<t},x,\pi^{\prime})}A^{t}\right]+\alpha\mathbb{E}\left[\sum_{t=1}^{\mathcal{T}}\Big{(}V(y_{<t},x,\pi)-\big{(}A_{t}+V(y_{<t},x,\pi^{\prime})\big{)}\Big{)}^{2}\right]$
By minimizing this loss, the policy model $\pi$ (the LLM under PPO training) is trained to generate explanations and refinements with higher reward but also constrained by not being distracted too much far away from the supervised-fine-tuned LLM $\pi^{\prime}$.
A.4 Additional results
A.4.1 Comparison with SFT on code generation
Table 9 shows the comparison between SFT CodeLlama-7B trained with our collected data (including the code generation, code explanation, and code refinement data), and the CodeLlama-7B trained with code generation data only. We find that although training with code generation enables LLMs to get comparable (or even higher pass@k on HumanEval) pass@k on one round of code generation, the LLMs cannot obtain strong self-debugging ability from code generation data. The LLM trained with code generation data can only improve the pass@k very little after self-debugging. By contrast, the LLM trained with the full collected data gets significantly higher pass@k after refinement, showing its strong self-debugging ability.
Table 9: Pass@k on MBPP and HumanEval by CodeLlama-7B trained on code generation only, and that trained with our collected code explanation and refinement data.
Models Approaches MBPP Humanval MBPP+ HumanEval+
pass@1 pass@10 pass@1 pass@10 pass@1 pass@10 pass@1 pass@10
Init. 46.18 70.56 40.14 70.37 41.23 60.58 34.22 63.34
SFT on code Refine 49.18 72.78 42.91 74.57 43.13 61.81 36.12 65.77
generation data Expl. + Refine 49.58 73.07 43.88 75.19 43.58 61.69 36.37 66.24
Init. 48.87 70.89 36.99 69.95 42.97 62.69 30.76 62.52
SFT on the Refine 58.07 77.34 52.65 80.71 51.64 71.04 46.61 74.43
full collected data Expl. + Refine 57.98 77.92 52.98 82.22 51.55 70.94 47.62 75.54
A.4.2 Iterative refinement
Figure 8 illustrates the pass@k metric when employing trained CodeLlama-13B for iterative code refinement. The findings align with those depicted in Figure 4, that both SFT and RL consistently outperform the prompting approach in each round of refinement. Notably, even after three rounds of refinement, the prompting method fails to achieve a higher pass@k than what SFT and RL models attain after just one round. This highlights the substantial advantage of SFT and RL methods in enhancing code quality. Moreover, the SFT and RL CodeLlama-13B models demonstrate robust continuous refinement capabilities, maintaining their superior performance across multiple iterations. This consistent outperformance underscores the effectiveness of SFT and RL strategies in refining and improving code, suggesting their potential for more efficient and reliable coding practices in iterative development scenarios.
Figure 8: Pass@k of prompting, SFT, and RL CodeLlama-13B after three iterations of refinements.
A.4.3 Generalizability
The pass@k of CodeLlama-13B trained with data collected from CodeLlama-34B are shown in Table 10, where RL achieves the highest pass@1 and pass@10 in all the benchmarks. The overall results are shown in Table 11, and SFT improves the pass@1 by up to 9.14% and pass@10 by 3.47% on the overall of MBPP and HumanEval, and improves the pass@1 by up to 8.68% and pass@10 by 3.92% on MBPP+ and HumanEval+. Further RL training improves the pass@1 by up to 0.61% and pass@10 by 0.86%. The improvement brought by RL is significantly larger on MBPP+ and HumanEval+ (up to 6.80% higher pass@1 and 2.28% higher pass@10), which is consistent with our finding when using GPT-3.5-Turbo’s training data that RL brings more considerable improvement on harder benchmarks.
Table 10: Pass@k on MBPP and HumanEval by LLMs trained with CodeLlama-34B’s data.
Models Approaches MBPP Humanval MBPP+ HumanEval+
pass@1 pass@10 pass@1 pass@10 pass@1 pass@10 pass@1 pass@10
Init. 42.88 70.85 37.11 74.69 38.93 62.26 30.15 66.27
Prompt. Refine 49.68 75.85 45.78 81.07 46.22 70.14 37.62 72.68
Expl. + Refine 49.97 76.39 45.90 81.18 45.77 70.48 38.36 73.84
Init. 50.61 73.14 44.96 76.85 44.66 63.15 37.98 69.59
CodeLlama-13B SFT Refine 59.07 79.52 54.15 83.95 52.87 73.50 46.96 77.40
Expl. + Refine 59.11 79.56 55.04 84.75 53.97 73.66 47.73 78.17
Init. 50.50 73.26 45.62 78.14 44.72 63.40 38.79 71.98
RL Refine 59.15 80.21 55.71 85.33 57.74 74.03 56.56 82.23
Expl. + Refine 59.57 80.18 56.08 85.40 56.60 73.17 56.24 82.39
Table 11: Overall pass@k on MBPP & HumanEval and MBPP+ & HumanEval+, trained with CodeLlama-34B’s data.
CodeLlama MBPP & HumanEval MBPP+ & HumanEval+
13B SFT RL SFT RL
Refine Expl. + Ref. Refine Expl. + Ref. Refine Expl. + Ref. Refine Expl. + Ref.
pass@1 57.85 (+9.13) 58.10 (+9.14) 58.30 (+0.45) 58.71 (+0.61) 50.46 (+7.74) 51.43 (+8.68) 57.26 (+6.80) 56.45 (+5.02)
pass@10 80.61 (+3.47) 80.84 (+3.27) 81.47 (+0.86) 81.47 (+0.63) 75.09 (+3.92) 75.5 (+3.65) 77.37 (+2.28) 76.92 (+1.42)
Table 12: Success refinement rate of different approaches over four benchmarks, trained with CodeLlama-34B’s data.
Models Refine (%) Explain + Refine (%)
Prompt. SFT RL Prompt. SFT RL
CodeLlama-13B 11.64 15.60 (+3.96) 22.74 (+7.14) 11.97 16.83 (+4.86) 21.61 (+4.78)
Table 12 shows the success refinement rate generated by CodeLlama-13B trained with data collected from CodeLlama-34B. The result is averaged over four benchmarks. SFT successfully refines 15.60 – 16.83% wrong solutions, outperforming the prompting approach by 3.96 – 4.86%. RL training further increases the refinement rate by 4.78 – 7.14% over SFT. We note that using CodeLlama-34B’s data, the improvement brought by SFT is smaller compared with training with GPT-3.5-Turbo’s data, which may be due to the slightly lower quality of code explanation and refinement data from CodeLlama-34B. Yet, RL training brings the refinement rate to a similar level to that trained with GPT-3.5-Turbo’s data, suggesting that RL training can potentially overcome the quality difference from open-sourced LLMs to commercial LLMs.
A.5 Case studies
A.5.1 Correct refinements generated by SFT LLMs
Figure 9 shows an example from the HumanEval benchmark for which the prompting approach fails to generate correct refinement. The bug in the initial solution is that ‘nlargest‘ returns the largest elements in descending order, but from the example provided, one can find out that the expected output is in ascending order. The prompting approach does not work for this example, and the prompted StarCoder generates a hallucinated explanation and simply repeats the wrong solution. Actually, such simple repeats of wrong solutions are very common when using the prompting approach, which supports our motivation to train LLM to self-debug. However, after SFT training, the trained StarCoder can correctly figure out the reason for test failure and generate the correct refinement.
Figure 9: Example for which prompting StarCoder fails to but SFT StarCoder generates the correct explanation and refinement.
A.5.2 Correct refinements only generated by RL LLMs
Figure 10 shows an example from the MBPP benchmark for which only the StarCoder trained with reinforcement learning explains the wrong code correctly, pointing out that the failing test case is because "this approach does not consider the case where multiple elements have the same maximum count" (highlighted with green background), and generates the correct refinement. The SFT StarCoder is unable to diagnose the wrong code correctly.
Figure 10: Example for which the RL LLM generates the correct explanation and refinement.
A.5.3 Robust refinements generated by RL LLMs
Figure 11 shows an example from HumanEval, for which both SFT and RL CodeLlama-13B generate the correct refinements that pass all the test cases from the HumanEval benchmark. Yet, the refinement from CodeLlama-13B is not fully correct, as “(x, y, z) == ((int)x, (int)y, (int)z)” is not equivalent to checking if these numbers are integers. The refinement generated by SFT CodeLLama-13B (left) fails the harder test cases in HumanEval+, i.e., AssertionError: expect any_int(3.0, 4, 7) == False, but was True, since it fails to realize that 3.0 is not integer. By contrast, the refinement generated by the RL CodeLlama-13B generates better refinement that passes all the test cases from HumanEval+.
Figure 11: Example for which RL LLM generates more robust refinement.
A.6 Human rating of code explanation
We let developers rate the explanations generated by different LLMs based on the following rubrics, which consider both the correctness of the statements in the explanations and also helpfulness of the explanation.
Table 13: Rubrics used for developer rating of LLMs’ generated explanations.
Rates 1 2 3 4 5
Correctness Totally wrong or misleading Mostly wrong but with at least one minor correct point Partially correct explanation with about 50% wrong Partially correct explanation with minor mistakes Totally correct explanation
Helpfulness Helpless statement that repeats the error or is not related to the bug Not very helpful but at least related to the bug Provides very vague hints about how to understand the failed test case Provides vague hints about how to fix the wrong code Provides at least one clear hints about how to fix
Figure 12: Example of human rating, where RL StarCoder generates the best explanation.
Figure 13: Example of human rating, where both SFT and RL StarCoder generate good explanations.
Figure 12 shows an example of developer rating, where the prompting and SFT StarCoder are rated “1” since their explanation is wrong and not informative. yet, the RL StarCoder’s explanation successfully points out that the function should return ‘Not Matched!’ when there is no match.
Figure 13 shows another example, where the prompting StarCoder is rated “2”, as the developer thinks the explanation is stating a correct fact. By contrast, SFT and RL StarCoder correctly point out that the function should check if the array is empty or not.
A.7 Potential impact
Using LLMs to help coding is popular nowadays. This work proposes a technique to train LLMs to explain and self-refine code, which aims to improve developers’ coding experience with LLMs. We also call for training LLMs to take feedback beyond the prompting approach to improve the LLMs’ self-debugging ability.
The technique is supposed to be on the same track as all the existing LLMs for code generation. Thus, we think no special concerns about broader impact need to be highlighted here.